# HeLo:::: Heterogeneous Multi-Modal Fusion with Label Correlation for Emotion Distribution Learning

### v2

@@ 1. Introduction
Emotion recognition plays a pivotal role in enhancing human-computer interaction (HCI) (Shneiderman and Plaisant, 2010; Li et al., 2023; Can et al., 2023), offering substantial benefits across various fields such as healthcare, education, and customer service (Alarcao and Fonseca, 2017). Researchers have investigated various behavioral signals, including facial expressions (Verma et al., 2021) and voice signals (Yi and Mak, 2020), as well as physiological signals like electroencephalogram (EEG) (Ye et al., 2022), electrocardiogram (ECG) (Hsu et al., 2017), and electromyogram (EMG) (Cheng and Liu, 2008), to detect emotional states. While many studies have concentrated on identifying basic emotions using the valence-arousal (VA) (Zhu et al., 2023) or valence-arousal-dominance (VAD) (Lang, 1995) model, there is an increasing awareness of the complexity and mixed nature of human emotions. Research indicates that individuals may simultaneously experience multiple emotions with varying intensities at the same time (Williams and Aaker, 2002; Zhao et al., 2020). However, most existing emotion recognition methods tend to focus on single-class models (Jia et al., 2019; Zhang et al., 2018; Zheng and Lu, 2016), which may not fully capture the varied, complex, and sometimes ambiguous nature of human emotions. Accurate identification of these complex emotional states is essential, as it reflects more realistic human experiences and interactions, underscoring the need for advanced emotion recognition systems that can interpret these nuances effectively.
@@
Researchers have explored the efficacy of integrating multi-modal data to enhance emotion recognition performance. Peng et al. (Peng et al., 2024) proposed CARAT to model fine-grained modality-to-label dependencies and exploit the modality complementarity by a shuffle-based aggregation strategy for multi-modal multi-label emotion recognition. Shu et al. (Shu et al., 2022) proposed an EDL convolutional neural network to predict the emotion distribution from the stacked physiological signals. Liu et al. (Liu et al., 2023) developed EmotionDict to learn a set of basic emotion elements in the latent feature space for emotion distribution learning. However, as the diagram shown in Figure 1 (a), these methods ignore the intrinsic correlation and heterogeneity among different modalities, resulting in a sub-optimal recognition performance. What’s more, they cannot fully exploit the label correlation across different basic emotions. As can be seen from Figure 1 (b), “afraid” and “scared” may be highly correlated. Leveraging the label correlations as a guidance can enhance the performance of emotion distribution learning.
@@ 2.2. Muli-modal Emotion Recognition
As multi-modal data provide a more comprehensive understanding of human emotional states (Ngai et al., 2022; Zheng et al., 2023), multi-modal emotion recognition has emerged as an important topic in the field of affective computing. The dominant multi-modal emotion recognition methods can be divided into two subfields: single-label emotion recognition (SLER) and multi-label emotion recognition (MLER). SLER recognized emotions using a single emotion label for each discrete time. MAET (Jiang et al., 2023) processed multi-modal inputs with specialized attention modules and alleviates subject discrepancy via adversarial training. While the goal of MLER is to assign an arbitrary number of emotion category labels to an input sample. Zhang et al. (Zhang et al., 2020) integrated transformer and multi-head modality attention to effectively fuse text, visual, and audio data and predict emotions. Compared with SLER and MLER, emotion distribution learning (EDL) has gained increasing attention, as it evaluates the intensity of each basic emotion, which is crucial for a comprehensive representation of complex emotional space. EmotionDict (Liu et al., 2023) utilized behavior and physiological responses to offer complementary information for emotion recognition and disentangled emotion features of an emotional state into a weighted combination of a set of basic emotion elements. Shu et al.(Shu et al., 2022) proposed EDLConv the combine multi-modal peripheral physiological signals for emotion distribution learning. In this paper, we propose a novel EDL framework for integrating multi-modal emotional data, and the model’s performance is improved with the guidance of learned label correlation.
@@ 3.2. Heterogeneity Mining with Optimal Transport
where $\mathcal{WD}(,)$ denotes the wasserstein distance, $\Pi\left(u,v\right)=\{T\in\mathbb{R}_{+}^{2C\times 2C}\mid T1_{2C}=u,T^{T}1_{2C}=v\}$, $1_{2C}$ represents a $2C$-dimensional all-one vector, $<,>_{F}$ refers to the Frobenius dot product, and $\text{Cost}\geq 0\in\mathbb{R}^{2C\times 2C}$ is a cost matrix calculated by the cosine distance between $x^{Phy}$ and $x^{v}$. The matrix $T\in\mathbb{R}^{2C\times 2C}$ is denoted as the optimal transport matching flow. Then, we apply the Pytorch framework-based batch-wise implementation of the Sinkhorn algorithm proposed in (Chen et al., 2020) to solve $\mathcal{WD}$. In this way, the obtained optimal transport matching flow $T$ can be used as a “cross-modal correlation” to enhance the multi-modal fusion. As such, the final multi-modal representation is adopted as follows:
@@ 3.3. Label Correlation-driven Cross-Attention
In the emotion distribution learning scenario, correlation across different labels plays a vital role in the prediction of emotional distributions. Therefore, we design a label correlation-driven cross-attention mechanism (LCDCA), shown in Figure 3, to effectively enhance the prediction performance of the model. Specifically, we introduce a learnable label embedding $x^{L}\in\mathbb{R}^{l\times d}$, where $l$ is the number of emotion classes. To learn the label correlation, we first generate learnable label correlation $M^{L}\in\mathbb{R}^{l\times l}$ and ground-truth label correlation $M^{gt}\in\mathbb{R}^{l\times l}$ through the cosine similarity, whose $(i,j)$-th element is given as:
@@
Table 1. Subject-dependent comparison of experimental results of our method and 10 comparison methods on six measures on the DMER and WESAD datasets. ↓ indicates “the smaller the better”, and ↑ indicates “the larger the better”. The parentheses show the corresponding ranks on each evaluation metric and the Average Ranks. The boldfaced scores denote the best performances and the underlined scores are the second runners. The results of the proposed HeLo are presented in a gray background.
@@ 4.3. Implementation Details
In our experiments, following previous works (Shu et al., 2022; Liu et al., 2023), both subject-dependent and subject-independent settings are adopted to verify the method on the two datasets. In subject-dependent recognition, the data of each subject are split into the training set and testing set in a ratio of 8:2 randomly, and the average performance across all subjects is taken as the final result. For subject-independent recognition, we utilize the leave-one-subject-out cross-validation for the evaluation, and the average performance across all iterations is reported as the final result. Our proposed framework is trained on a 12GB NVIDIA GeForce RTX3080 GPU and developed under PyTorch. The training adopts Adam optimizer with a learning rate of $10^{-3}$. The batch size is set to 128, the number of epochs for training is set to 300, and the number of attention heads is 4. The embedding size in the attention mechanism and feed-forward network is set to 128, and 64, respectively, and we set the depth of transformer blocks to 1. Our code will be released soon.
@@ 4.4. Comparison with State-of-the-art Methods
Subject-Dependent Chebyshev (↓) 0.0529 0.0574 0.0487 0.0552 0.0446
@@
Subject-Independent Chebyshev (↓) 0.0977 0.0953 0.0964 0.0926 0.0882
@@
Subject-Dependent Chebyshev (↓) 0.0088 0.0074 0.0079 0.0082 0.0073
@@
Subject-Independent Chebyshev (↓) 0.0502 0.0419 0.0429 0.0474 0.0403
@@
Table 2 illustrates the subject-independent evaluation of the comparison methods and our proposed model on the DMER and WESAD datasets. Except for the KL metric on the WESAD dataset, our method achieves the best results across both datasets and shows the highest average rank. The subject-independent recognition is more challenging than the subject-dependent situation, showing the superiority of our proposed method in capturing the subject-invariant emotional features across different individuals.
@@
Figure 5. Ablation of network components in the proposed method. The experiment is conducted on the DMER and WESAD datasets, using both subject-dependent and subject-independent evaluation.
@@ 4.5. Ablation Studies
Ablation on the modalities. We also assessed the impact of each modality by separately removing each modality from our full model, the results are shown in Table 3 and Table 4. These experiments were conducted in both subject-dependent and subject-independent settings across both datasets. The results indicate that each modality contributes significantly to the overall performance of multi-modal emotion distribution learning, effectively integrating multiple modalities can enhance the emotion distribution performance. More ablation results on modalities can be referred to the supplementary materials.
Figure 6. Effect of the number of attention heads. It shows the performance over six metrics on the DMER dataset.
Figure 7. The effect of the depth of transformer block in the model on six evaluation metrics on the DMER dataset.
Ablation on the number of attention heads. The number of attention heads is usually an essential parameter for models based on attention mechanisms. Thus, we also compared the effect of different choices of the number of heads on the six evaluation metrics of the model on the DMER dataset in both subject-dependent and subject-independent settings, shown in Figure 6. It can be seen that the model’s performance first increases across subject-dependent and subject-independent settings as the number of attention heads increases. Then, the model achieves the best performance across all six evaluation metrics when the number of attention heads is equal to 4, after which the performance declines as the number of heads continues to increase. However, the number of attention heads has a relatively limited impact on the performance of the model.
Ablation on the depth of transformer block. The depth of end-to-end models is considered a crucial factor in determining their fitting ability. Therefore, we evaluate the model while varying the depth of the block on the DMER dataset to explore the impact of the depth of the transformer block. Figure 7 shows the optimal classification performance was achieved when the depth of the block was 1, and as the depth increased, the classification accuracy tended to slightly decrease and leveled off as the depth continued to increase, as the greater the depth of the transformer block, the more likely it is to overfit the model.
@@ 4.6. Visualization
Visualization of the cross-modal representations. We visualize the multi-modal feature space in the testing set by utilizing t-SNE (Van der Maaten and Hinton, 2008), and the results are illustrated in Figure 8. In Figure 8 (a), before the OTHM module, there is a noticeable separation between physiological (blue) and behavioral (red) representations. This pronounced heterogeneity illustrates the challenges faced in multi-modal fusion due to distinct feature distributions, which can impede effective integration and analysis. With the help of OTHM, shown in Figure 8 (b), the cross-modal representations exhibit a closer arrangement and overlap more significantly than before. This reduction in heterogeneity suggests that OTHM has effectively minimized the disparities between the physiological and behavioral features, facilitating a better-aligned feature space and a more coherent multi-modal fusion.
@@
Conformer MAET CARAT EmotionDict
FLOPs 2.422M 91.955M 330.671M 156.354M 11.656M
Params 1.023M 15.672M 6.780M 0.483M 4.040M
4.7. Model Efficiency
@@ 5. Conclusion
In this paper, we introduce a novel multi-modal emotion distribution learning framework, HeLo, which effectively addresses the challenges of capturing complex, mixed emotional states from heterogeneous physiological and behavioral data. The CAPF module ensures the robust fusion of multi-modal physiological signals, and the OTHM module efficiently mines the heterogeneity between physiological and behavioral representations, bridging the gap between these diverse data sources and enhancing feature alignment. Moreover, the LCDCA fully exploits label correlations, guiding the learning of emotion distributions in a semantically meaningful way. Experimental results on two multi-modal emotion datasets demonstrate the efficacy of our proposed method. In the future, we plan to utilize LLMs to extract the label embeddings, which provide strong semantic priors to our model instead of random initialization.
@@ Appendix D More Ablation Results on Modalities
We conducted thorough ablation experiments on different combinations of different emotion modalities. The results are illustrated in Table 9, 10, 11, and 12. The experimental results demonstrate the necessity and effectiveness of our proposed multi-modal emotion distribution learning framework.

### v3

# HeLo::: Heterogeneous Multi-Modal Fusion with Label Correlation for Emotion Distribution Learning

@@ 1. Introduction
Emotion recognition plays a pivotal role in enhancing human-computer interaction (HCI) (Shneiderman and Plaisant, 2010; Li et al., 2023; Can et al., 2023), offering substantial benefits across various fields such as healthcare, education, and customer service (Alarcao and Fonseca, 2017). Researchers have investigated various behavioral signals, including facial expressions (Verma et al., 2021) and voice signals (Yi and Mak, 2020), as well as physiological signals like electroencephalogram (EEG) (Ye et al., 2022), electrocardiogram (ECG) (Hsu et al., 2017), and electromyogram (EMG) (Cheng and Liu, 2008), to detect emotional states. While many studies have concentrated on identifying basic emotions using the valence-arousal (VA) (Zhu et al., 2024) or valence-arousal-dominance (VAD) (Lang, 1995) model, there is an increasing awareness of the complexity and mixed nature of human emotions. Research indicates that individuals may simultaneously experience multiple emotions with varying intensities at the same time (Williams and Aaker, 2002; Zhao et al., 2020). However, most existing emotion recognition methods tend to focus on single-class models (Jia et al., 2019; Zhang et al., 2018; Zheng and Lu, 2016), which may not fully capture the varied, complex, and sometimes ambiguous nature of human emotions. Accurate identification of these complex emotional states is essential, as it reflects more realistic human experiences and interactions, underscoring the need for advanced emotion recognition systems that can interpret these nuances effectively.
@@
Researchers have explored the efficacy of integrating multi-modal data to enhance emotion recognition performance (Liu et al., 2025). Peng et al. (Peng et al., 2024) proposed CARAT to model fine-grained modality-to-label dependencies and exploit the modality complementarity by a shuffle-based aggregation strategy for multi-modal multi-label emotion recognition. Shu et al. (Shu et al., 2022) proposed an EDL convolutional neural network to predict the emotion distribution from the stacked physiological signals. Liu et al. (Liu et al., 2023) developed EmotionDict to learn a set of basic emotion elements in the latent feature space for emotion distribution learning. However, as the diagram shown in Figure 1 (a), these methods ignore the intrinsic correlation and heterogeneity among different modalities, resulting in a sub-optimal recognition performance. What’s more, they cannot fully exploit the label correlation across different basic emotions. As can be seen from Figure 1 (b), “afraid” and “scared” may be highly correlated. Leveraging the label correlations as a guidance can enhance the performance of emotion distribution learning.
@@ 2.2. Muli-modal Emotion Recognition
As multi-modal data provide a more comprehensive understanding of human emotional states (Ngai et al., 2022; Zheng et al., 2023; Zhu et al., 2025), multi-modal emotion recognition has emerged as an important topic in the field of affective computing. The dominant multi-modal emotion recognition methods can be divided into two subfields: single-label emotion recognition (SLER) and multi-label emotion recognition (MLER). SLER recognized emotions using a single emotion label for each discrete time. MAET (Jiang et al., 2023) processed multi-modal inputs with specialized attention modules and alleviates subject discrepancy via adversarial training. While the goal of MLER is to assign an arbitrary number of emotion category labels to an input sample. Zhang et al. (Zhang et al., 2020) integrated transformer and multi-head modality attention to effectively fuse text, visual, and audio data and predict emotions. Compared with SLER and MLER, emotion distribution learning (EDL) has gained increasing attention, as it evaluates the intensity of each basic emotion, which is crucial for a comprehensive representation of complex emotional space. EmotionDict (Liu et al., 2023) utilized behavior and physiological responses to offer complementary information for emotion recognition and disentangled emotion features of an emotional state into a weighted combination of a set of basic emotion elements. Shu et al.(Shu et al., 2022) proposed EDLConv the combine multi-modal peripheral physiological signals for emotion distribution learning. In this paper, we propose a novel EDL framework for integrating multi-modal emotional data, and the model’s performance is improved with the guidance of learned label correlation.
@@ 3.2. Heterogeneity Mining with Optimal Transport
[…]
@@ 3.3. Label Correlation-driven Cross-Attention
[…]
@@
Table 1. Subject-dependent comparison of experimental results of our method and 10 comparison methods on six measures on the DMER and WESAD datasets. ↓ indicates “the smaller the better”, and ↑ indicates “the larger the better”. The parentheses show the corresponding ranks on each evaluation metric and the Average Ranks. The boldfaced scores denote the best performances and the underlined scores are the second runners. The results of the proposed HeLo are presented in a blue background.
@@ 4.3. Implementation Details
In our experiments, following previous works (Shu et al., 2022; Liu et al., 2023), both subject-dependent and subject-independent settings are adopted to verify the method on the two datasets. In subject-dependent recognition, the data of each subject are split into the training set and testing set in a ratio of 8:2 randomly, and the average performance across all subjects is taken as the final result. For subject-independent recognition, we utilize the leave-one-subject-out cross-validation for the evaluation, and the average performance across all iterations is reported as the final result. Our proposed framework is trained on a 12GB NVIDIA GeForce RTX3080 GPU and developed under PyTorch. The training adopts Adam optimizer with a learning rate of $10^{-3}$. The batch size is set to 128, the number of epochs for training is set to 300, and the number of attention heads is 4. The embedding size in the attention mechanism and feed-forward network is set to 128, and 64, respectively, and we set the depth of transformer blocks to 1. Our code is released at: https://github.com/kaio-99/HeLo.
@@ 4.4. Comparison with State-of-the-art Methods
Subject-Dependent
Chebyshev (↓) 0.0529 0.0574 0.0487 0.0552 0.0446
@@
Subject-Independent
Chebyshev (↓) 0.0977 0.0953 0.0964 0.0926 0.0882
@@
Subject-Dependent
Chebyshev (↓) 0.0088 0.0074 0.0079 0.0082 0.0073
@@
Subject-Independent
Chebyshev (↓) 0.0502 0.0419 0.0429 0.0474 0.0403
@@
Table 2 illustrates the subject-independent evaluation of the comparison methods and our proposed model on the DMER datasets. Except for the KL metric on the WESAD dataset, our method achieves the best results across both datasets and shows the highest average rank. The subject-independent recognition is more challenging than the subject-dependent situation, showing the superiority of our proposed method in capturing the subject-invariant emotional features across different individuals.
@@
Figure 5. Ablation of network components in HeLo.
@@ 4.5. Ablation Studies
Ablation on the modalities. We also assessed the impact of each modality by separately removing each modality from our full model, the results are shown in Table 3 and 4. These experiments were conducted in both subject-dependent and subject-independent settings across both datasets. The results indicate that each modality contributes significantly to the overall performance of multi-modal emotion distribution learning, effectively integrating multiple modalities can enhance the emotion distribution performance.
Figure 6. Effect of attention heads on the DMER
Figure 7. Effect of transformer block depth on the DMER.
Ablation on the number of attention heads. The number of attention heads is usually an essential parameter for models based on attention mechanisms. The result in Figure 6 indicates that, the model’s performance in both subject-dependent and subject-independent settings initially improves as the number of heads increases, and subsequently declines. The model reaches its optimal performance across all six evaluation metrics when the number of attention heads was 4. However, the overall impact of the number of attention heads on the model’s performance is relatively limited.
Ablation on the depth of transformer block. The depth of end-to-end models is considered a crucial factor in determining their fitting ability. Figure 7 illustrates that, the model achieved optimal classification performance when the transformer block depth was 1. As the depth increased, the classification accuracy showed a slight tendency to decrease and then leveled off. This is because a greater transformer block depth might cause overfitting.
@@ 4.6. Visualization
Visualization of the cross-modal representations. We visualize the multi-modal feature space in the testing set by utilizing t-SNE (Van der Maaten and Hinton, 2008). In Figure 8 (a), before the OTHM module, there is a noticeable separation between physiological (blue) and behavioral (red) representations. This pronounced heterogeneity illustrates the challenges faced in multi-modal fusion due to distinct feature distributions, which can impede effective integration and analysis. With the help of OTHM, shown in Figure 8 (b), the cross-modal representations exhibit a closer arrangement and overlap more significantly than before. This reduction in heterogeneity suggests that OTHM has effectively minimized the disparities between the physiological and behavioral features, facilitating a better-aligned feature space and a more coherent multi-modal fusion.
@@ 5. Conclusion
In this paper, we introduce a novel multi-modal emotion distribution learning framework, HeLo, which effectively addresses the challenges of capturing complex, mixed emotional states from heterogeneous physiological and behavioral data. The CAPF module ensures the robust fusion of multi-modal physiological signals, and the OTHM module efficiently mines the heterogeneity between physiological and behavioral representations, bridging the gap between these diverse data sources and enhancing feature alignment. Moreover, the LCDCA fully exploits label correlations, guiding the learning of emotion distributions in a semantically meaningful way. Experimental results on two multi-modal emotion datasets demonstrate the efficacy of our proposed method.
@@ Appendix D More Ablation Results on Modalities
We conducted thorough ablation experiments on different combinations of different emotion modalities. The results are illustrated in Table 8, 9, 10, and 11. The experimental results demonstrate the necessity and effectiveness of our proposed multi-modal emotion distribution learning framework.
Conformer MAET CARAT EmotionDict
FLOPs 2.422M 91.955M 330.671M 156.354M 11.656M
Params 1.023M 15.672M 6.780M 0.483M 4.040M
Appendix E Model Efficiency